What is the width of rectangle (w=3x-7) (l=2x+3) if the perimeter is 132 cm?
I have the answer but not sure how to work the problem... Thank you.

Answer by lynnlo(4176)

FINDING THE WIDTH OF A RECTANGLE,WITH THE GIVEN PERIMETER
P IS PERIMETER
L IS LENGTH
W IS WIDTH
P=2L+2W
P=2(L+W)
P/2=L+W
OK, NOW IF THE PERIMETER IS GIVEN YOU GOT ONE SIDE OF THE
MEASURE LENGTH(THEN YOU CAN FIND THE MEASURE LENGTH OF THE OTHER SIDE
;;;;;;;IF YOU KNOW THE MEASURE LENGTH THEN THE MEASURE LENGTH OF
W(WIDTH) IS P/2-L=W AND VICE VERSA